< 0) = 1/3, and Exercise 9.8. Suppose X has an N(u,02) distribution, P(X P(X < 1) = 2/3. What are the values of u and o?!
Answer #1

let X~ N(M, 82) Given - P(X50)=1/3 & P(XS1)=2/3 .. P(x(0)=1/3 P(x(1) = 2/3 Peru 0-) = 0.33 PO AY < 1% = 0.67 .P(24-0.44) = 0.